OpenStudy (anonymous):

Solve for x: 5/x^2-1 - 2/x=2/x+1? A. x= -1/4 and x= -2 B. x= 1/4 and x= -2 C. x= -1/4 and x=2 D. x= 1/4 and x=2

OpenStudy (anonymous):

\[\frac{ 5 }{x^2 }-1-\frac{2}{x}=\frac{2}{x}+1 => \frac{5}{x^2}-\frac{-4}{x}-2=0\] you could simply plug in values for x and see which ones work, or rewrite the equation and solve for x using quadratic formula